A steady current of 36 mA exists in a wire. How many electrons pass any given point in the wire per second? electrons per second
Answer #1

Steady current means an electric current that does not change with time.

so current flow in a wire at any time,ampere law, i=\frac{N\Delta q}{\Delta t}

Here i=36mA=36\times10-3Ampere

time,t=1second

Charge of electron ,q=1.6\times10-19coulomb

so number of electrons pass per second,=\frac{36\times 10^{-3}}{1.6\times 10^{-19}}=22.5\times 10^{16} electrons